20191219-Optics-Mtg

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> HRS_Optics_Mtg


previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 PARTICIPANT CODE:                  #382 744 247
 Room IRL:                            <none>


Agenda

Minutes

  1. KP recaps goals: get Q2 and FOM, figure out inelastic contribution (which enters into FOM calculation by selection of dp cutoff)
  2. Siyu reports that he has finished fitting Dec 8 optics data and has new data bases posted (by Ye Tian) on the analysis cluster.
    1. He is called in from this hospital, but don't worry, everything is fine
    2. These fits only incorporate dp=0, but use all sieve holes
    3. Nilanga will discuss these results with him tomorrow to make sure the new db should be kosher
    4. After discussion, we conclude that the Dec 8 data was taken in the current tune "E" (Elmer Fudd).
    5. Siyu will also use new (sept 15) optics data to test his database.
    6. We should update all Q2 measurements etc with the new optics DB to be sure we have correct calculations
  3. Devi haplog entry haplog:3979
    1. shows alignment spectra for Ca-48 and C-12.
    2. There is a shift between elastic peaks, explicable by combination of dE/dx and nuclear recoil
  4. Ryan shows preliminary FOM work (not yet posted)
    1. now extending calculation to average over acceptance (rather than taking a central value)
    2. Some discussion of the Q2. Apparently the left Q2 is reported 4% below the right Q2.
      1. not matched by rate difference
      2. may be resolved by updated DB. We'll see, soon.
    3. Ryan also showed his simulation results for Q2 (0.0297 GeV2), which were about 1% below the right Q2 (0.030 GeV2) and 3% above the left Q2 (0.0289 GeV2)
  5. Dustin asks if we can get an estimate of the asymmetry in the out-of-plane acceptance beween the left and right arms from sieve data
    1. Siyu has a macro that can provide a cut for any given sieve hole. He will share it.
  6. Next steps: (This wasn't explicitly discussed but is my best understanding)
    1. Ryan will continue his FOM study. He will also update his data by re-running with the new data base.
    2. Siyu will examine the Ca-48 sieve data to try to fit elastic peak strengths
    3. Devi will post Ca-48 alignment, sieve in, and Q2 runs to optics run-list google sheet CREX Optics Runlist
    4. Someone needs to post Q2 results updated with the new data base. It wasn't clear to me who would do it, but I'm desperate to see it so you know I'm gonna ask someone.
    5. We will post results to haplog and communicate via slack over the break
  7. Next meeting, same time, 2 weeks away (1pm on Thursday, Jan 2, 2020)

Attendence

Bob, Devi, Dustin, Siyu, Ryan, Nilanga, Kent, Sanghwa